OpenStudy (anonymous):

What is the range of the function f(x) = 3x2 + 6x – 8

OpenStudy (richie63):

lemme check

OpenStudy (richie63):

Manipulate your equation into the "vertex form": y= a(x-h)^2+k where (h,k) is the vertex . f(x) = 3x^2 + 6x - 7 f(x) = (3x^2 + 6x) - 7 f(x) = 3(x^2 + 2x) - 7 Complete the square: f(x) = 3(x^2 + 2x + 1) - 7 - 3 f(x) = 3(x+1)^2 - 10 . Therefore, the vertex is: (h,k) = (-1, -10) . To find the y-intercepts, set x=0 and solve for f(x): f(x) = 3x^2 + 6x - 7 f(x) = 3(0^2) + 6(0) - 7 f(x) = - 7 y-intercept is (0, -7) . To find x-intercepts, set f(x)=0 and solve for x: 0 = 3x^2 + 6x - 7

OpenStudy (richie63):

answer at the bottom bruh

OpenStudy (anonymous):

ight thanks

OpenStudy (xmissalycatx):

You're not supposed to give direct answers @Richie63 . I understand the work and all but you should have just backspaced on the actual answer b/c you can get into trouble for that.. And I already know you got the answer from a website ;)